OpenStudy (anonymous):

Mixing Problem... I know I have derived the correct equation from the paragraph. I just need help with the differentiation.

OpenStudy (anonymous):

\[\frac{ dx }{ dt }=.4-\frac{ 2x }{ 25 }\rightarrow \frac{ 10-2x }{ 25 }\]

OpenStudy (anonymous):

I need it in the form x(t)

OpenStudy (anonymous):

Im thinking separation of variables

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ dx }{ 10-2x }=\int\limits_{}^{}\frac{ dt }{ 25 }\]

OpenStudy (anonymous):

\[-\frac{ 1 }{ 2 }\ln (10-2x)=\frac{ t }{ 25 }+C\]

OpenStudy (anonymous):

If you see a mistake stop me or else I'm going to keep going

hartnn (hartnn):

thats correct, go on...

OpenStudy (anonymous):

x(0) = 0.005

OpenStudy (anonymous):

based on the problem.

OpenStudy (anonymous):

Now this is where i get tripped up. Do I have to get it in the form x(t)=, or can I just plug .005 and 0 in right now?

hartnn (hartnn):

u can plug it right now...

hartnn (hartnn):

u will get C

OpenStudy (anonymous):

ok C = ln 9.99

hartnn (hartnn):

nopes.

OpenStudy (anonymous):

I need to find t when x(t)=.02

OpenStudy (anonymous):

ok let me try again

OpenStudy (anonymous):

-(ln 9.99)/2

OpenStudy (anonymous):

the first time I multiplied the right side by -2 and assumed C could just absorb the -2 because it is a constant times a constant. That is breaking a rule?

hartnn (hartnn):

now u are correct and yes.

OpenStudy (anonymous):

So now I just simply plug in .02 for x and solve for t

hartnn (hartnn):

exactly!

OpenStudy (anonymous):

The back of the book says (0.4)(100-t)-(3.9E-7)(100-t)^4; t=19.96mins. 100L is the volume of the tank, .4 is the initial concentration. But I dont see how they got the rest of their equation

OpenStudy (anonymous):

I do feel like we did it right

OpenStudy (anonymous):

in flow and out flow are the same

hartnn (hartnn):

if there's 100-t^4 then there is some thing missing in your original equation....if your original equation is correct, then we are correct.

OpenStudy (anonymous):

in flow rate is 8L/min and is equal to the out flow. The volume of the tank is kept constant at 100L. Initially there is 0.5kg of salt in the tank. Concentration of salt entering the tank is 0.05kg/L.

OpenStudy (anonymous):

So what I did was dy/dx= Conc(initial)*InFlow rate - outFlow rate*x(t) over 100L

OpenStudy (anonymous):

where x(t) is the kg of salt. So when divide by the Volume gives me the final concentration.

hartnn (hartnn):

i donno how to form the equation with your given word problem.....but if your first equation is correct, then all your next math part is correct...

OpenStudy (anonymous):

I copied the setup from a very similar problem in the book so I think my first equation is right. The book went from \[\frac{ dy }{ dx }=0.6-\frac{ 3x }{ 500 }\rightarrow x(t)=100(1-e ^{-3t/500)}\]This said it gives them the weight of the salt then they divided the equation by the volume to get the equation for the concentration of salt at time t. I forgot to do that part, but my integration doesn't look like there's

OpenStudy (anonymous):

Thanks for your help. If you want to pull the rip cord on this question and help others I'd understand

hartnn (hartnn):

thanks and sorry....

OpenStudy (anonymous):

my example above is from a different but similar problem

OpenStudy (anonymous):

no big